Lista dei presenti alla prova scritta del 16 aprile 2007 Matricola

10045476  Insufficiente
10042734  Insufficiente
10036911  Ammesso/a all’orale
10043090  Insufficiente
10048374  Ammesso/a all’orale
10028352  Ammesso/a all’orale
10044252  Ammesso/a all’orale
10042441  Insufficiente
10044524  Insufficiente
10019925  Insufficiente
10041688  Insufficiente
10028497 o 10028477  Insufficiente